# Zero-Shot Temporal Resolution Domain Adaptation for Spiking Neural Networks

### v2

@@ 3.3.1 Event-based data
We first provide preliminaries for the notion of temporal resolution for event-based data. Event-based data can be represented as asynchronous or synchronous. In the asynchronous representation, data is provided as a list of events with each their precise time of occurrence, while in the synchronous representation, data is provided in a time-binned format as spike sequences of a given length, each bin being populated with all events occurring during this period of time. The first type of representation is naturally adopted when data comes from event-based sensors, such as NMNIST [orchard2015converting] and the second type of data representation is typically used after post-processing of data obtained by conventional sensors, such as Spiking Heidelberg Digits (SHD) dataset [shddataset], or by post-processing of data from event-based sensors.
The standard approach for processing event-based data with SNNs is to use digital clock-based SNNs, which require the data to be in an effectively synchronous form for processing. Hence, if the data is asynchronous, it is converted into a synchronous form, such as by binning events on regular time intervals.
@@ 3.3.2 Temporal resolution
In the below, we base our exposition on synchronous data representation. Consider a continuous-time signal $\bm{s}(t)$, $0\leq t\leq t_{c}$ with a duration $t_{c}$ seconds, representing a phenomena of interest, such as a spoken keyword. We denote by $\bm{s}\in\mathbb{Z}^{N\times 1}$ an integer-valued discrete-time sequence extrapolated from $\bm{s}(t)$ where $N$ is the number of time steps. Hence, for a regular discretization of time-axis, which we focus on here, a single time step in $\bm{s}$ will represent a time duration of $\Delta_{N}=t_{c}/N$ seconds.
Suppose that we have two discrete-time data sequences associated with the same continuous signal, each with a different number of time steps. Let us refer to them as $\bm{s}_{S}\in\mathbb{Z}^{N_{S}\times 1}$ and $\bm{s}_{T}\in\mathbb{Z}^{N_{T}\times 1}$ with $N_{S}\neq N_{T}$, and thus $\Delta_{S}\neq\Delta_{T}$. Hence, $\bm{s}_{S}$ and $\bm{s}_{T}$ correspond to sequences with different time resolutions. The sequences $\bm{s}_{S}$ and $\bm{s}_{T}$ can be obtained from $\bm{s}(t)$ in various ways. One is to obtain both of the sequences directly from $\bm{s}(t)$ using different encoding parameters. Another is to obtain $\bm{s}_{S}$ directly from $\bm{s}(t)$, while obtaining $\bm{s}_{T}$ by transforming $\bm{s}_{S}$, or vice versa.
@@ 3.3.3 Model adaptation
We consider the scenario where the SNN model is pre-trained on data with time resolution $\Delta_{S}$ but we would like to use this model on data with temporal resolution $\Delta_{T}$, where $\Delta_{T}\neq\Delta_{S}$. Accordingly, we refer to training dataset as source data, and the later as target data. Similarly, we refer to $\Delta_{S}$ and $\Delta_{T}$ as temporal resolution of source and target data, respectively.
During the SNN training, two groups of parameters are learned for each neuron, parameters of the neuron dynamic, $\bm{H}_{v,(S)},\bm{H}_{f,(S)},\bm{H}_{i,(S)},\bm{H}_{r,(S)}$, and the weights of the synaptic connections between the neurons, $\bm{W},\bm{V}$. Our main goal in this article is to adjust the model parameters such that we obtain high-accuracy results on target data. We propose to achieve this goal by keeping the dynamics of the state variable $\bm{v}$ of neurons, see (2), as similar as possible under the source and target data. This means the state variable $\bm{v}$ dynamics is preserved even when the temporal resolution of the data changes.
We consider this problem under two constraints: First, we are only allowed to adapt the neuron dynamic parameters, $\bm{H}_{k}$ for $k=v,i,f,r$. Hence, SNN architecture and synaptic connections are unchanged. Second, the only known information about the target data is the ratio between the target and source temporal resolution. Hence, we do not rely on any post-deployment re-training on the target dataset, which can be costly and prone to interference. Our results in Section LABEL:sec:numericalResults show that promising performance can be obtained even under this constrained scenario, supporting the central role of neuron dynamics for capturing time-dependent behaviour.
To summarize, the main goal of this article is to find a mapping from the source to target feature space in the form of
@@
where the mapping ${\cal{M}}(.)$ represents the temporal resolution domain adaptation, and the parameter $\rho$ is given by
@@
  • Coarse-to-Fine deployment: The model is pretrained on source data with coarse temporal resolution and tested on target data with fine temporal resolution, i.e., $\Delta_{S}>\Delta_{T}$, thus $\rho<1$.
  • Fine-to-Coarse deployment: The model is pretrained on source data with fine temporal resolution and tested on target data with coarse temporal resolution, i.e., $\Delta_{S}<\Delta_{T}$, thus $\rho>1$.
@@ 4.1 Preliminaries - Correspondence between SNNs and non-linear SSMs
A general discrete-time nonlinear SSM can be written as [OkuyamaYoshifumi2014DCS]:
@@
where $\bm{v}[t]\in\mathbb{R}^{n\times 1},\bm{f}[t]\in\mathbb{R}^{r\times 1}$, $\bm{y}[t]\in\mathbb{R}^{p\times 1}$ represent the state vector, the input to the system and the output of the system, respectively, and $\mathcal{F}(\cdot)$ and $\mathcal{G}(\cdot)$ are possibly non-linear functions. Here $n$ is the dimension of the system state, $r$ is dimension of the system input, and $p$ is dimension of the system output. The general neuron model in (2) can be expressed as a non-linear SSM as in (6) using
$\!\mathcal{F}(\bm{v}[t],\bm{f}[t])$ $=\bm{H}_{v}\bm{v}[t]+\bm{H}_{f}g_{\bf{\Theta}}(\bm{v}[t])+\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}\bm{f}[t]$ (7a)
@@
where $\mathcal{F}_{A}(\cdot),\mathcal{F}_{B}(\cdot),\mathcal{G}_{C}(\cdot)$ and $\mathcal{G}_{D}(\cdot)$ are possibly non-linear functions. The generalized neuron (2) can be expressed as a non-linear SSM with separable functions (8) using
@@
$\mathcal{F}_{B}(\bm{f}[t])$ $=\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}\bm{f}[t]$ (9b)
@@ 4.2 Preliminaries - Correspondence between SNNs and linear SSMs
As linear SSMs enable analytical investigations and form the basis of our adaptation methods, in this section, we propose an approximate correspondence between the general SNN neuron model in (2) and linear SSMs. Despite SNNs being strictly non-linear SSMs, this approximate correspondence can be utilized to achieve promising results for temporal domain adaptation, as demonstrated by the results in Section LABEL:sec:numericalResults.
A linear SSM in discrete-time can be written as [Gajic]:
@@
To find an approximate correspondence between the general neuron model in (2) and (10), we first map the nonlinear function in the separable non-linear SSM (8) to the matrices describing linear relation in (10). The mapping is intuitive,
@@
where the symbol $\rightarrow$ indicates that the left-hand side linear term is best described as part of the dynamics of the right-hand side possibly nonlinear function. Next, we associate the general neuron matrices $\bm{H}_{v},\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}$ in (2), to the linear matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ in (10). Using the correspondence in (9) and (11), the following mapping arises as a candidate: $\bm{A}=\begin{bmatrix}\bm{H}_{v},\bm{H}_{f}\end{bmatrix}^{T}$, and $\bm{B}=\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}$. However, we propose to make a deviation from this mapping and interpret the spike of the previous time step as external input to the neuron state. Hence, we use $\bm{A}=\bm{H}_{v}$, and $\bm{B}=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}$. To visualize this correspondance, we rewrite the general spiking neuron in (2) as
$\!\bm{v}[t+1]$ $=\bm{H}_{v}\bm{v}[t]+\begin{bmatrix}\bm{H}_{f}\\ \bm{H}_{i}\\ \bm{H}_{r}\end{bmatrix}\begin{bmatrix}s_{out}[t]\\ \bm{W}\bm{s}_{in}[t]\\ \bm{V}\bm{s}_{out}[t]\end{bmatrix}$ (12a)
@@
$\bm{v}[t+1]$ $=\bm{A}\bm{v}[t]+\bm{B}\bm{f}[t]$ (13a)
where $\bm{v}\in\mathbb{R}^{n\times 1},\bm{A}=\bm{H}_{v}\in\mathbb{R}^{n\times n},\bm{B}=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}\in\mathbb{R}^{n\times r}$, and $\bm{f}[t]=\begin{bmatrix}s_{out}[t],\bm{W}\bm{s}_{in}[t],\bm{V}\bm{s}_{out}[t]\end{bmatrix}^{T}\in\mathbb{R}^{r\times 1}$ such that $\bm{f}$ combines all input spikes to the neuron including the effect of its own output. Since only the $\bm{A}$ and $\bm{B}$ matrices are needed for adapting the state variable dynamics, which is our main interest here, we leave (9c) and (9d) as they are, since these functions are associated with the output spikes $\bm{y}[t]$ and are not required for adaptation of the state variable.
@@ 4.3 Proposed Methods for Temporal Resolution Domain Adaptation
In this section, we define our proposed temporal resolution domain adaptation methods for the generalized SNN neuron model in (2). As explained in Section 3.3, the proposed methods suggest a mapping ${\cal{M}}(.)$ from the source domain with temporal resolution $\Delta_{S}$ to the target domain with temporal resolution $\Delta_{T}$, where $\rho=\frac{\Delta_{T}}{\Delta_{S}}$.
Note that all the methods presented below have been derived in detail in the form of propositions in Section LABEL:sec:propositions.
Integral approximation adaptation method is derived from the Integral approximation method for obtaining discrete-time SSM from a continuous-time system model. Following Proposition LABEL:Integral_Method_Prop in Section LABEL:sec:propositions, the proposed mapping is
@@
Euler adaptation method is derived from the Euler approximation method for obtaining discrete-time SSM from a continuous-time system model. Following Proposition LABEL:Euler_Method_Prop in Section LABEL:sec:propositions, the proposed mapping is
@@
Expectation adaptation method is based on the assumption the expected values of the state variable vector between different time resolutions are the same. This adaptation method assumes that the ratio between the time steps ($\Delta$) of the lower and higher resolutions is an integer. Specifically, if $\rho>1,\rho\in\mathbb{Z}_{+}$, while if $\rho<1,1/\rho\in\mathbb{Z}_{+}$. Following Proposition LABEL:Method_Expectation_Prop in Section LABEL:sec:propositions the proposed mapping is
@@
$\!\bm{H}_{k,(T)}$ =$=\begin{cases}(\sum^{\rho}_{j=1}\left(\bm{H}_{v,(S)}\right)^{\rho-j})\bm{H}_{k,(S)}&\rho\geq 1\\ (\sum^{1/\rho}_{j=1}\left(\bm{H}_{v,(T)}\right)^{1/\rho-j})^{-1}\bm{H}_{k,(S)}&\rho<1\end{cases}$ (16b)
@@
In the Integral and Expectation adaptation methods, $\left(\bm{H}_{v,(S)}\right)^{\rho}$ represents a matrix raised to a possibly fractional power which can be computed using the Schur–Padé algorithm [MatrixFractionalPower]. In general, fractional powers of a real-valued matrix may be complex-valued. By restricting the parameters of the adLIF neuron (1) to the ranges $\alpha,\beta,a\in(0,1)$, the possibly fractional power $\rho=1/\bar{\rho}$ of the real matrix $\bm{H}_{v,(S)}$ used in our experiments remains real-valued [MatrixFractionalPower].

### v3

@@ 3.3.1 Event-based data
In this section, we state the event-based data representation we focus on in this article. Event-based data can be represented as asynchronous or synchronous [20, 8]. In the asynchronous representation, data is provided as a list of events each with their precise time of occurrence. In the synchronous representation, data is provided in a time-binned format as spike sequences of a given length, each bin populated with all events occurring during this period of time. The first type of representation is naturally adopted when data comes from event-based sensors, such as NMNIST [45] and the second type of data representation is typically used after post-processing of data obtained by conventional sensors, such as Spiking Heidelberg Digits (SHD) dataset [12], or by post-processing of data from event-based sensors.
The standard approach for processing event-based data with SNNs is to use digital clock-based SNNs, which require the data to be in an effectively synchronous form for processing. Hence, if the data is asynchronous, it is converted into a synchronous form, such as by binning events on regular time intervals. In this article, we focus on such synchronous data representations where events that occur during a temporal window of fixed length are accumulated to create sequences of spikes [20, 37]. This type of representations have been used in a large number of earlier SNN works, including studies utilizing temporal bins for audio signals [6, 24] and frame-based event representation for spatio-temporal data [zubić2024state, 7, 25].
@@ 3.3.2 Temporal resolution
In this section, we define the notion of temporal resolution used in this article. Consider a continuous-time signal $\bm{s}(t)$, $0\leq t\leq t_{c}$ with a duration $t_{c}$ seconds, representing a phenomenon of interest, such as a spoken keyword. We denote by $\bm{s}\in\mathbb{Z}^{N\times 1}$ an integer-valued discrete-time sequence extrapolated from $\bm{s}(t)$ where $N$ is the number of time steps. Hence, for a regular discretization of the time-axis, which we focus on here, a single time step in $\bm{s}$ will represent a time duration of $\Delta_{N}=t_{c}/N$ seconds.
Suppose that we have two discrete-time data sequences associated with the same continuous signal, each with a different number of time steps. Let us refer to them as $\bm{s}_{S}\in\mathbb{Z}^{N_{S}\times 1}$ and $\bm{s}_{T}\in\mathbb{Z}^{N_{T}\times 1}$ with $N_{S}\neq N_{T}$, and thus $\Delta_{S}\neq\Delta_{T}$. Hence, $\bm{s}_{S}$ and $\bm{s}_{T}$ correspond to sequences with different time resolutions. The sequences $\bm{s}_{S}$ and $\bm{s}_{T}$ can be obtained from $\bm{s}(t)$ in various ways. One is to obtain both sequences directly from $\bm{s}(t)$ using different encoding parameters. Another is to obtain $\bm{s}_{S}$ directly from $\bm{s}(t)$, while obtaining $\bm{s}_{T}$ by transforming $\bm{s}_{S}$, or vice versa.
@@ 3.3.3 Model adaptation
We consider the scenario where the SNN model is pre-trained on data with time resolution $\Delta_{S}$ but we would like to use this model on data with temporal resolution $\Delta_{T}$, where $\Delta_{T}\neq\Delta_{S}$. Consequently, we refer to the training dataset as source data and the latter as target data. Similarly, we refer to $\Delta_{S}$ and $\Delta_{T}$ as the temporal resolution of the source and target data, respectively.
During SNN training, two groups of parameters are learned: parameters of the neuron dynamics for each neuron, $\bm{H}_{v,(S)},\bm{H}_{f,(S)},\bm{H}_{i,(S)},\bm{H}_{r,(S)}$, and the weights of the synaptic connections between the neurons, $\bm{W},\bm{V}$. Our main goal in this article is to adjust the model parameters such that we obtain high-accuracy results on target data. We propose to achieve this goal by keeping the dynamics of the state variable $\bm{v}$ of neurons, see (2), as similar as possible under the source and target data. This means that the dynamics of the state variable $\bm{v}$ is preserved even when the temporal resolution of the data changes.
We consider this problem under two constraints: First, we are only allowed to adapt the neuron’s dynamic parameters, $\bm{H}_{k}$ for $k=v,i,f,r$. Hence, SNN architecture and synaptic connections are unchanged. Second, the only known information about the target data is the ratio between the target and the source temporal resolution. Hence, we do not rely on any post-deployment re-training on the target dataset, which can be costly. Our results in Section 5 show that promising performance can be obtained even under this constrained scenario, supporting the central role of neuron dynamics in capturing time-dependent behavior.
In summary, the main goal of this article is to find a mapping from the source to the target feature space in the form of
@@
[…]
@@
  • Coarse-to-Fine deployment: The model is pre-trained on source data with coarse temporal resolution and tested on target data with fine temporal resolution, i.e., $\Delta_{S}>\Delta_{T}$, thus $\rho<1$.
  • Fine-to-Coarse deployment: The model is pre-trained on source data with fine temporal resolution and tested on target data with coarse temporal resolution, i.e., $\Delta_{S}<\Delta_{T}$, thus $\rho>1$.
@@ 4.1 Preliminaries - Correspondence between SNNs and non-linear SSMs
A general discrete-time nonlinear SSM can be written as [44]:
@@
where $\bm{v}[t]\in\mathbb{R}^{n\times 1},\bm{f}[t]\in\mathbb{R}^{r\times 1}$, $\bm{y}[t]\in\mathbb{R}^{p\times 1}$ represent the state vector, the input to the system and the output of the system, respectively, and $\mathcal{F}(\cdot)$ and $\mathcal{G}(\cdot)$ are possibly non-linear functions. Here, $n$ is the dimension of the system’s state, $r$ is the dimension of the system’s input, and $p$ is the dimension of the system’s output. The general neuron model in (2) can be expressed as a non-linear SSM as in (7) using
$\!\mathcal{F}(\bm{v}[t],\bm{f}[t])$ $=\bm{H}_{v}\bm{v}[t]+\bm{H}_{f}g_{\bf{\Theta}}(\bm{v}[t])+\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}\bm{f}[t]$ (8a)
@@
where $\mathcal{F}_{A}(\cdot),\mathcal{F}_{B}(\cdot),\mathcal{G}_{C}(\cdot)$ and $\mathcal{G}_{D}(\cdot)$ are possibly non-linear functions. The generalized neuron in (2) can be expressed as a non-linear SSM with separable functions as in (9) using
@@
$\mathcal{F}_{B}(\bm{f}[t])$ $=\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}\bm{f}[t]$ (10b)
@@ 4.2 Preliminaries - Correspondence between SNNs and linear SSMs
As linear SSMs enable analytical investigations and form the basis of our adaptation methods, in this section, we propose an approximate correspondence between the general SNN neuron model in (2) and linear SSMs. Despite SNNs being strictly non-linear SSMs, this approximate correspondence can be utilized to achieve promising results for temporal domain adaptation, as demonstrated by the results in Section 5.
A linear SSM in discrete-time can be written as [19]:
@@
To find an approximate correspondence between the general neuron model in (2) and the linear SSM in (11), we first map the nonlinear function in the separable non-linear SSM (9) to the matrices describing linear relation in (11). The mapping is intuitive,
@@
where the symbol $\rightarrow$ indicates that the left-hand side linear term is best described as part of the dynamics of the right-hand side possibly nonlinear function. Next, we associate the general neuron matrices $\bm{H}_{v},\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}$ in (2), to the linear matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ in (11). Using the correspondences in (10) and (12), the following mapping arises as a candidate: $\bm{A}=\bm{H}_{v}+\bm{H}_{f}\bm{G}_{\bf{\Theta}}$ and $\bm{B}=\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}$, where $\bm{G}_{\bf{\Theta}}\in\mathbb{R}^{1\times n}$ represents a linearization of the non-linear spiking function $g_{\bf{\Theta}}(\cdot)$. However, we propose to make a deviation from this mapping and interpret the spike of the neuron itself as an external input to the neuron state. This deviation avoids the linearization of the non-linear spiking function. Hence, we use $\bm{A}=\bm{H}_{v}$, and $\bm{B}=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}$. To visualize this correspondence, we rewrite the general spiking neuron in (2) as
$\!\bm{v}[t+1]$ $=\bm{H}_{v}\bm{v}[t]+\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}\begin{bmatrix}s_{out}[t]\\ \bm{W}\bm{s}_{in}[t]\\ \bm{V}\bm{s}_{out}[t]\end{bmatrix}$ (13a)
@@
$\bm{v}[t+1]$ $=\bm{A}\bm{v}[t]+\bm{B}\bm{f}[t]$ (14)
where $\bm{v}\in\mathbb{R}^{n\times 1},\bm{A}=\bm{H}_{v}\in\mathbb{R}^{n\times n},\bm{B}=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}\in\mathbb{R}^{n\times r}$, and $\bm{f}[t]=\begin{bmatrix}s_{out}[t],\bm{W}\bm{s}_{in}[t],\bm{V}\bm{s}_{out}[t]\end{bmatrix}^{T}\in\mathbb{R}^{r\times 1}$ such that $\bm{f}$ combines all input spikes to the neuron, including the effect of its own output. Since only the $\bm{A}$ and $\bm{B}$ matrices are needed for adapting the state variable dynamics, which is our main interest here, we leave (10c) and (10d) as they are, since these functions are associated with the output spikes $\bm{y}[t]$ and are not required for adaptation of state dynamics.
We now consider adLIF neuron as an example. For adLIF, using (3.2), we obtain the following as $\bm{A}$ and $\bm{B}$ of (14):
$\bm{A}$ $=\bm{H}_{v}=\begin{bmatrix}\alpha&-(1-\alpha)\\ a&\beta\end{bmatrix}$, (15a)
$\bm{B}$ $=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}=\begin{bmatrix}-\alpha\theta&1-\alpha&1-\alpha\\ b&0&0\end{bmatrix}$. (15b)
@@ 4.3 Proposed Methods for Temporal Resolution Domain Adaptation
[…]
Note that all the methods presented below have been derived in detail in the form of propositions in Section A.1. A step-by-step pseudo-code outlining the usage of the methods is provided in Algorithm 1.
Integral approximation adaptation method is derived from the Integral approximation method for obtaining discrete-time SSM from a continuous-time system model. Following Proposition 27 in Section A.1, the proposed mapping is
@@
Euler adaptation method is derived from the Euler approximation method for obtaining discrete-time SSM from a continuous-time system model. Following Proposition 29 in Section A.1, the proposed mapping is
@@
Expectation adaptation method is based on the assumption the expected values of the state variable vector between different time resolutions are the same. This adaptation method assumes that the ratio of the time steps ($\Delta$) of the lower and higher resolutions is an integer. Specifically, if $\rho>1,\rho\in\mathbb{Z}_{+}$, while if $\rho<1,1/\rho\in\mathbb{Z}_{+}$. Following Proposition 3 in Section A.1 the proposed mapping is
@@
$\!\bm{H}_{k,(T)}$ $=\begin{cases}(\sum^{\rho}_{j=1}\left(\bm{H}_{v,(S)}\right)^{\rho-j})\bm{H}_{k,(S)}&\rho\geq 1\\ (\sum^{1/\rho}_{j=1}\left(\bm{H}_{v,(T)}\right)^{1/\rho-j})^{-1}\bm{H}_{k,(S)}&\rho<1\end{cases}$ (18b)
@@
In the Integral and Expectation adaptation methods, $\left(\bm{H}_{v,(S)}\right)^{\rho}$ represents a matrix raised to a possibly fractional power which can be computed using the Schur–Padé algorithm [27]. In general, fractional powers of a real-valued matrix may be complex-valued. By restricting the parameters of the adLIF neuron (1) to the ranges $\alpha,\beta,a\in(0,1)$, the possibly fractional power $\rho=1/\bar{\rho}$ of the real matrix $\bm{H}_{v,(S)}$ used in our experiments remains real-valued [27].
(1) Train on source resolution
Input : Training dataset $\mathcal{D}_{S}$ with source resolution $\Delta_{S}$.
Initialize model randomly $SNN_{S}=(\bm{H}_{v,(S)}^{(0)},\bm{H}_{i,(S)}^{(0)},\bm{H}_{f,(S)}^{(0)},\bm{H}_{r,(S)}^{(0)},\bm{W}_{S}^{(0)},\bm{V}_{S}^{(0)})$.
for epoch $1,\dots,E$ do
  Update the $SNN_{S}$ model parameters using $\mathcal{D}_{S}$.
return $SNN_{S}=(\bm{H}_{v,(S)},\bm{H}_{i,(S)},\bm{H}_{f,(S)},\bm{H}_{r,(S)},\bm{W}_{S},\bm{V}_{S})$.
1ex(2) Perform adaptation
Input : Trained $SNN_{S}$, $\rho$, i.e., the ratio $\frac{\Delta_{T}}{\Delta_{S}}$, choice of adaptation method, i.e., Integral, Euler or Expectation.
Adapt neuron model parameters of $SNN_{S}$ i.e. set $\bm{H}_{v,(T)}\bm{H}_{i,(T)},\bm{H}_{f,(T)},\bm{H}_{r,(T)}={\cal{M}}(\rho,\bm{H}_{v,(S)},\bm{H}_{i,(S)},\bm{H}_{f,(S)},\bm{H}_{r,(S)})$ using (16) for Integral (17) for Euler, or (18) for Expectation.
Create the model $SNN_{T}=(\bm{H}_{v,(T)},\bm{H}_{i,(T)},\bm{H}_{f,(T)},\bm{H}_{r,(T)},\bm{W}_{S},\bm{V}_{S})$.
return $SNN_{T}$.
1ex(3) Use on data with target resolution
Input : $SNN_{T}$, testing dataset $\mathcal{D}_{T}$ with target resolution $\Delta_{T}$.
Output : Performance of $SNN_{T}$ on $\mathcal{D}_{T}$ i.e., $SNN_{T}(\mathcal{D}_{T})$.
Algorithm 1 Pseudo-code for the work-flow with the proposed adaptation methods